### v1

# Motion Before Action: Diffusing Object Motion as Manipulation Condition

Abstract
Inferring object motion representations from observations enhances the performance of robotic manipulation tasks. This paper introduces a new paradigm for robot imitation learning that generates action sequences by reasoning about object motion from visual observations. We propose MBA, a novel module that employs two cascaded diffusion processes for object motion generation and robot action generation under object motion guidance. MBA first predicts the future pose sequence of the object based on observations, then uses this sequence as a condition to guide robot action generation. Designed as a plug-and-play component, MBA can be flexibly integrated into existing robotic manipulation policies with diffusion action heads. Extensive experiments in both simulated and real-world environments demonstrate that our approach substantially improves the performance of existing policies across a wide range of manipulation tasks.
Index Terms:
Robotic Manipulation, Imitation Learning
I Introduction
Physiological research shows that humans process complex object motion information in their environment to support effective action execution [Human]. This motion analysis enables an understanding of object dynamics, which in turn guides human actions such as reaching, grasping, and maneuvering around obstacles.
In contrast to the biological approach, most existing robot policies [ACT, octo, DP, RISE, cage, DP3] are predominantly guided by observation, employing feature encoder and adopting generative approaches to predict actions. While effective, it often results in an overreliance on environmental cues, with the model focusing on memorizing observation features rather than reasoning about object motion patterns, as humans do. Consequently, when encountering extensive pose shifts in real-world objects or actions, many policies often struggle to generalize effectively [octo, openvla], which can limit their practical performance.
To address these challenges and improve execution capabilities, we aim to equip the robot with human-like reasoning skills by inferring future object motion from observations, and then predicting future actions under the object motion guidance. By achieving these objectives, the robot can derive intrinsic information (such as poses and motions) from the scene, allowing the policy to map observations to actions in a way that aligns more closely with human reasoning, i.e., reasoning about object motion rather than simply memorizing actions [jannerreasoning].
In this work, we introduce Motion Before Action (referred to as MBA), a novel module that equips existing robotic manipulation policies with such reasoning skills. We observe that representing the object motion using 6D pose aligns with the robot’s end-effector pose in the same format, making the actions compatible with this representation. Consequently, they exhibit mathematical consistency. During task execution, at each time step, the robot pose, object pose, and action are in proximity within the same space, with their kinematic relationships following a learnable pattern, demonstrating physical consistency [xu2021end]. Based on these two distributional consistencies, we hypothesize the learnability of object pose representation in diffusion models [DDPM, DDIM]: 1) diffusion models have been shown to effectively reconstruct actions [DP], so they should similarly excel at reconstructing analogous object poses; 2) just as the robot’s pose can serve as a condition [octo] for generating actions via diffusion models, so does the object pose.
In this paper, we make the following key contributions:
  • We propose a novel imitation learning paradigm for robotic manipulation that allows robots to extract object pose sequences from observations and use them to aid in action prediction, thereby enhancing the robustness and kinematic consistency of the policy’s observation-to-action mapping.
  • We introduce MBA, a flexible auxiliary module that can be easily integrated into existing policies with diffusion heads, rather than functioning as an independent policy.
  • We conduct comparative experiments with MBA on three 2D and 3D robotic manipulation policies with diffusion action heads [DP, DP3, RISE], demonstrating substantial performance improvements across various tasks. These tasks, comprising 57 tasks from 3 simulation benchmarks and 3 real-world tasks, involve articulated object manipulation, soft and rigid body manipulation, tool use, non-tool use, and diverse action patterns. Results show that MBA consistently enhances the performance of such policies in both simulated and real-world environments.
II Related Work
II-A Imitation Learning for Robotic Manipulation
Imitation learning seeks to enable robots to acquire complex skills by learning from expert demonstrations. Recently, behavior cloning [bc] from demonstrations has demonstrated promising performance across various robotic manipulation tasks [BCZ, rt1, rt2, oxe, octo, openvla]. Most existing robotic manipulation policies are typically structured into three main components: an observation perception module for manipulation-relevant information extraction from observations, an optional policy backbone for processing the perception results, and an action head for generating actions based on the (processed) perception outputs.
For action generation, prior work has investigated both single- and multi-step action prediction methods. Single-step action prediction [rt1, rt2, oxe, openvla, zhang2024leveraging] is straightforward but often leads to inconsistencies between the consecutive actions. To address this limitation, several studies have applied the action chunking technique [ACT] to enable multi-step action sequence prediction. This includes the utilization of unimodal action heads, such as L1 and L2 action heads [ACT, roboagent, liu2024foam, kareer2024egomimic], and multimodal action heads like the Gaussian mixture head [MandlekarWhat, bet, vqbet, cuiplay] and diffusion head [octo, DP, RISE, cage, DP3]. Among these, the diffusion head [DDPM] excels in capturing the diversity and complexity of action sequences, which is essential for executing real-world tasks.
Similar to the multimodal nature of robot action sequences in manipulation tasks, object motion sequences also exhibit diverse and complex characteristics, making diffusion heads well-suited for modeling such dynamics. Accordingly, we develop the MBA module, designed for seamless integration into policies with diffusion action heads [DP, DP3, RISE]. This module enhances performance by decoupling the process into two stages: generating object motion via a motion diffusion head and conditionally generating robot actions through an action diffusion head.
II-B Object-Centric Manipulation Learning
Object-centric manipulation learning can be categorized into two main lines of approaches: one emphasizing scene understanding through structured, object-centric representations [groot, viola, orion, tyree20226, migimatsu2020object, devin2018deep], and the other focusing on action-oriented learning by identifying how objects can be manipulated to achieve specific goals [bahl2023affordances, sriramahrp, XuFlowCrossDomain, BharadhwajTrack2act, YuanGeneralFlow].
The action-oriented approach actively engages with objects to achieve desired outcomes. Affordance learning is a central method here, aiming to identify possible actions based on an object’s physical properties and context [bahl2023affordances, SAM2_policy, BorjaAffordance, nasiriany2024rt, yuan2024robopoint, sriramahrp]. This approach, though effective in structured scenarios, often relies on predefined affordances, limiting its adaptability to dynamic environments and tasks that require fine-grained coordination of actions or adaptive responses.
To overcome these limitations, inspired by the success of flow-based methods [teed2020raft, goyal2022ifor], recent work has utilized object flow as a general affordance for guiding manipulation policies [WenAnyPoint, VecerikRobotap, XuFlowCrossDomain, BharadhwajTrack2act, YuanGeneralFlow, EisnerFlowbot3d, ZhangFlowbotPlusPlus, SeitaToolflownet]. These policies map vision to motion by anchoring visual features and constructing object flow, which is then translated into robot actions using either heuristic transformations [YuanGeneralFlow, BharadhwajTrack2act, ZhangFlowbotPlusPlus, SeitaToolflownet, EisnerFlowbot3d] or learned models [WenAnyPoint, XuFlowCrossDomain]. Many methods incorporate object pose as an intermediate representation [YuanGeneralFlow, BharadhwajTrack2act, SeitaToolflownet] in such flow-to-action translations. However, using flow as an indirect motion representation can introduce ambiguity due to the vision-motion gap [WenAnyPoint, XuFlowCrossDomain, BharadhwajTrack2act], complicating flow-to-action learning. Moreover, defining such heuristic transformations is challenging, often hindering generalization in dynamic or novel environments.
Our proposed MBA module employs object pose as a direct motion representation, which reduces training demands and enhances interpretability by establishing a clear link between visual inputs and motion outcomes. While a contemporaneous work [SPOT] similarly leverages object pose for motion representation, it infers robot actions heuristically from object trajectories. We argue that generating actions from object motion guidance is essential, particularly in tasks involving deformable objects or dynamic conditions where heuristic methods may fall short.
III Method
We aim to endow the policy with the capability to concurrently reason about object pose $\mathbf{M}$ and robot action $\mathbf{A}$ from observations $\mathbf{O}$. Specifically, we model the joint conditional distribution $p(\mathbf{M},\mathbf{A}|\mathbf{O})$ in MBA. This distribution can be decoupled as:
$p(\mathbf{M},\mathbf{A}|\mathbf{O})=p(\mathbf{M}|\mathbf{O})p(\mathbf{A}|\mathbf{M},\mathbf{O})$
where $p(\mathbf{M}|\mathbf{O})$ serves to sample the most likely object motions based on current observations, and $p(\mathbf{A}|\mathbf{M},\mathbf{O})$ serves to sample the most likely actions, guided by the sampled motions and current observations.
Figure 1: Overview of MBA. MBA takes the current observation as input, which could be in the form of 3D point clouds or RGB images from different viewpoints. Object pose sequences are sampled as actions with denoising diffusion from the object policy to be part of the framework’s output. Conditioning on the observations and object pose actions, MBA samples deployable robot actions with denoising diffusion from the robot policy. These actions are executed within the workspace to update the environment state and the observations.
MBA is an object motion generation module $p(\mathbf{M}|\mathbf{O})$ that is readily compatible with any existing policy with diffusion action heads, enhancing versatility and efficiency across various robotic manipulation contexts. Given an observation, the perception module of the original policy first processes it into observation features. These observation features are then passed to MBA, where they serve as a condition for the object pose sequence denoising process. Subsequently, the diffusion action head of the original policy takes both the observation features and the denoised object pose sequence as inputs and outputs the corresponding robot actions. During training, both ground-truth object pose sequences and action sequences are used as supervision for the predicted object pose sequences from MBA and action sequences from the action head. At inference time, the policy with MBA operates in an end-to-end manner, without requiring ground-truth object poses. The full execution framework of MBA is illustrated in Fig. 1.
III-A Object Motion Generation
As discussed in §II-B, we opt for object pose as a direct representation to describe object motion. The object pose $\mathrm{m}_{t}$ is represented as a 9D vector, combining 3D translation and 6D rotation [rot6d], to align with robot action representations [DP, DP3, RISE]. At each observation step $t$, we use the observation features $O_{t}$ as the condition to the diffusion process [DDPM], which denoises the object pose sequence for the next $T_{m}$ steps $\mathbf{M}_{t:t+T_{m}}=(\mathrm{m}_{t},\mathrm{m}_{t+1},\mathrm{m}_{t+2}...\mathrm{m}_{t+T_{m}})$.
Specifically, the diffusion-based object pose generation model begins by sampling an initial noise from a Gaussian distribution $\mathbf{M}^{k}\sim\mathcal{N}(0,I)$. At each diffusion step $k$, the denoising network $\varepsilon_{\theta}$ progressively removes this noise, conditioned on the observation features $O_{t}$. This process iterates to recover a noise-free, clean object pose sequence $\mathbf{M}^{0}$:
$\mathbf{M}^{k-1}=\alpha_{k}\left(\mathbf{M}^{k}-\gamma_{k}\varepsilon_{\phi}(\mathbf{M}^{k},O,k)\right)+\sigma_{k}\mathcal{N}(0,I),$ (1)
where $\{\alpha_{k},\gamma_{k},\sigma_{k}\}$ is the noise schedule.
To train the noise prediction network $\varepsilon_{\theta}$, we randomly pick a sequence of real object trajectory $\mathbf{M}^{0}$ and apply noise at a random diffusion step $k$ through a forward diffusion process. The model is then tasked to predict the added noise $\varepsilon^{k}$ at the corresponding diffusion step. We employ mean squared error (MSE) loss as the objective function to supervise the prediction of object pose:
$\mathcal{L}=\text{MSE}\left(\epsilon^{k},\varepsilon_{\phi}(\mathbf{M}^{0}+\epsilon^{k},{O},k)\right),$ (2)
where the ground-truth object pose $\mathbf{M}^{0}$ is obtained via a motion capture (MoCap) system during demonstration collection, and $\epsilon^{k}$ is the added noises for the diffusion step $k$.
III-B Robot Action Generation under Object Motion Guidance
Following [DP, DP3, RISE], the robot action $\mathrm{a}_{t}$ is defined by the end-effector pose and gripper width, using a 10D representation that concatenates 3D translation, 6D rotation [rot6d], and 1D gripper width together. Similarly, we use the diffusion process to generate the action sequence at the next $T_{a}$ steps: $\mathbf{A}_{t:t+T_{a}}=(\mathrm{a}_{t},\mathrm{a}_{t+1},\mathrm{a}_{t+2}...\mathrm{a}_{t+T_{a}})$.
In this process, we generate actions $\mathbf{A}^{0}$ using the same diffusion method as for object poses. Notably, the action noise prediction network receives not only the same observation features $O$ as before, but also the object pose features $M$ obtained in the previous stage. These features are concatenated to predict the noise jointly:
$\mathbf{A}^{k-1}=\alpha_{k}\left(\mathbf{A}^{k}-\gamma_{k}\varepsilon_{\phi}(\mathbf{A}^{k},M,O,k)\right)+\sigma_{k}\mathcal{N}(0,I)$ (3)
Given the physical and mathematical similarities between object poses and the robot’s end-effector poses, we use a similar encoding network for their feature representation. The object pose is unfolded along the $T_{m}$ axis into $T_{m}\times 9$ and then passed through MLP layers with dimensions $(T_{m}\times 9,32,32)$, ultimately resulting in the object pose feature vector $M$.
From every iteration, we collect an action sequence $\mathbf{A}^{0}$ from expert demonstrations and add noise over $k$ steps. The noise prediction network is then tasked with predicting the noise based on the feature information, with supervision provided through MSE loss:
$\mathcal{L}=\text{MSE}\left(\epsilon^{k},\varepsilon_{\varphi}(\mathbf{A}^{0}+\epsilon^{k},{M},{O},k)\right)$ (4)
III-C Execution
During execution, the policy with MBA estimates the object pose for $T_{m}$ steps based on the observation inputs, which include environmental information and the robot state. This estimated pose sequence serves as a direct condition for generating an action sequence of $T_{a}$ steps, without reliance on the MoCap system. The robot then executes up to $T_{a}$ action steps before the next observation and execution cycle, with the exact number of steps depending on task type and environmental conditions. To adhere to standard kinematic principles, we ensure that $T_{m}\geq T_{a}$ in all experiments. The specific process is detailed in Alg. 1.
Algorithm 1 MBA Execution
0:  Observation $\mathbf{O}_{t}$, Predetermined steps $T_{m},T_{a}^{\prime}\leq T_{a}$, Encoders $\mathbf{F}_{\mathrm{O}},\mathbf{F}_{\mathrm{M}}$
0:  Task completion via iterative action execution
  while task is not complete do
     Encode the observation: $O_{t}\leftarrow\mathbf{F}_{\mathrm{O}}(\mathbf{O}_{t})$
     Predict object pose sequence $\{\mathrm{m}_{t},\mathrm{m}_{t+1},\dots,\mathrm{m}_{t+T_{m}}\}$:
     for $k=K$ downto $1$ do
        $\mathbf{M}^{k-1}=\alpha_{k}\left(O^{k}-\gamma_{k}\epsilon_{\phi}(\mathbf{M}^{k},O_{t},k)\right)+\sigma_{k}\mathcal{N}(0,I)$
     end for
     $M_{t}\leftarrow\mathbf{F}_{\mathrm{M}}\left(\mathbf{M}^{0}=\{\mathrm{m}_{t},\mathrm{m}_{t+1},\dots,\mathrm{m}_{t+T_{m}}\}\right)$
     Predict action sequence $\{\mathrm{a}_{t},\mathrm{a}_{t+1},\dots,\mathrm{a}_{t+T_{a}}\}$ with extracted feature $M_{t}$ and $O_{t}$:
     for $k=K$ downto $1$ do
        $\mathbf{A}^{k-1}=\alpha_{k}\left(A^{k}-\gamma_{k}\epsilon_{\varphi}(\mathbf{A}^{k},M_{t},O_{t},k)\right)+\sigma_{k}\mathcal{N}(0,I)$
     end for
     Execute $T_{a}^{\prime}$ steps, $T_{a}^{\prime}\leq T_{a}$
     for $i=1$ to $T_{a}^{\prime}$ do
        Execute $\mathrm{a}_{t+i}$: $\mathrm{a}_{t+i}\leftarrow\begin{cases}\mathrm{a}_{t+i}&\text{If task condition holds}\\ \text{Stop}&\text{Otherwise}\end{cases}$
     end for
     Update observations $\mathbf{O}_{t}$
     $t\leftarrow t+T_{a}^{\prime}$
  end while
IV Simulation Experiments
In the simulation experiments, we aim to address the following research questions: (Q1) Can MBA effectively improve the performance of robotic manipulation policies by leveraging predicted object motion as a condition for generating robot actions? (Q2) Can the human-like reasoning process of MBA make policy learning more efficient?
Method Adroit (3) DexArt (4) MetaWorld
Easy
MetaWorld Medium (11) MetaWorld
Hard
MetaWorld VeryHard (5) Average (57)
DP3 $68.3\pm 3.3$ $53.5\pm 2.5$ $90.9\pm 1.4$ $61.6\pm 6.5$ $29.7\pm 2.8$ $49.0\pm 6.8$ $71.3\pm 3.2$
DP3 w. MBA $79.7\pm 0.7$ $52.3\pm 2.8$ $92.5\pm 1.1$ $66.4\pm 6.1$ $36.2\pm 1.2$ $86.8\pm 1.6$ $77.5\pm 2.2$
DP $31.7\pm 3.0$ $22.8\pm 3.8$ $83.6\pm 3.6$ $31.1\pm 5.3$ $9.0\pm 1.0$ $26.6\pm 3.2$ $53.6\pm 3.6$
DP w. MBA $64.0\pm 3.0$ $29.0\pm 2.5$ $84.9\pm 1.2$ $57.8\pm 4.4$ $23.8\pm 1.3$ $79.8\pm 1.2$ $67.8\pm 2.0$
TABLE I: Average success rate and standard deviation of policies integrating MBA compared with baselines under 3 random seeds for 57 simulation tasks.
IV-A Setup
Benchmarks. We evaluate our policy in three simulation benchmarks, encompassing a total of 57 tasks:
  • Adroit [Adroit] utilizes a multi-fingered Shadow robot in the MuJoCo [mujoco] environment to perform highly dexterous manipulation across a variety of tasks. These operations involve both articulated objects and rigid bodies.
  • DexArt [dexart] uses the Allegro robot in the SAPIEN [sapien] environment to perform high-precision dexterous manipulation. It primarily focuses on tasks involving articulated object manipulation.
  • MetaWorld [metaworld] primarily operates in the MuJoCo environment, using a gripper to perform manipulation tasks involving both articulated and rigid objects. It covers a wide range of skills required for everyday scenarios, categorizing these tasks into difficulty levels: easy, medium, hard, and very hard.
Baselines. The focus of this work is to demonstrate that the introduction of MBA as a module can universally enhance the performance of existing policies with diffusion heads. Therefore, in our simulation experiments, we select representative Diffusion Policy [DP] and 3D Diffusion Policy [DP3] as our 2D and 3D baseline respectively. We then integrate the MBA module into these baselines and compare the performance. To ensure fairness in the experiments, we ensure that MBA and the corresponding baseline methods use the same expert demonstrations for training, with an equal amount of training steps. During the execution phase, both methods undergo the same number of observation and inference steps.
Demonstrations. In terms of expert demonstration generation, we employ scripted policies for MetaWorld, the VRL3 [vrl3] agent for Adroit, and the PPO [ppo] agent for DexArt. The average demonstration success rates for these agents are 98.7%, 72.8%, and nearly 100%, respectively. For Adroit and MetaWorld, we use 10 demonstrations for training, while 100 demonstrations are used for DexArt.
Protocols. Following [DP3], We conduct 3 runs for each experiment, using seed numbers 0, 1, and 2. For each seed, we evaluate 20 episodes every 200 training epochs and compute the average of the top 5 success rates. We then report the mean and standard deviation of success rates across the 3 seeds.
IV-B Results
Method Adroit DexArt MetaWorld
Door Pen Laptop Toilet Bin Picking Box Close Hammer Peg Insert Side Disassemble Shelfplace Reach
DP3 $62\pm 4$ $43\pm 6$ $81\pm 2$ $71\pm 3$ $34\pm 30$ $42\pm 3$ $76\pm 4$ $69\pm 7$ $69\pm 4$ $17\pm 10$ $24\pm 1$
DP3 w. MBA $74\pm 1$ $65\pm 1$ $78\pm 6$ $70\pm 2$ $54\pm 23$ $56\pm 2$ $98\pm 2$ $75\pm 5$ $98\pm 1$ $73\pm 1$ $32\pm 5$
DP $37\pm 2$ $13\pm 2$ $31\pm 4$ $26\pm 8$ $15\pm 4$ $30\pm 5$ $15\pm 6$ $34\pm 7$ $43\pm 7$ $11\pm 3$ $18\pm 2$
DP w. MBA $45\pm 1$ $53\pm 3$ $42\pm 2$ $45\pm 2$ $45\pm 8$ $36\pm 7$ $89\pm 5$ $53\pm 0$ $71\pm 23$ $64\pm 2$ $26\pm 3$
TABLE II: Detailed performance of policies integrating MBA and baselines in several typical simulation tasks.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 2: The relationship between the success rates of MBA and baselines and the number of training steps in a single experiment.
Integrating MBA results in more stable and superior performance (Q1). We report the average success rates and standard deviations (across three random seeds) for all simulation tasks in the Adroit, Dexart, and MetaWorld environments, across four difficulty levels, in Tab. I. On average, MBA outperforms the baseline in the majority of benchmarks, achieving a 14.2% increase in the average success rate over DP and a 6.2% increase over DP3. Additionally, we observe a general reduction in the average standard deviation of task execution across all benchmarks, further confirming the robustness of MBA. The results also demonstrate that MBA significantly improves success rates in tasks with higher difficulty levels. Detailed examples are provided in Tab. II, where we report the average success rate and standard deviation for particular tasks. Notably, in tasks that require precise contact and fine manipulation within narrow action spaces, the accurate object pose estimation and motion prediction powered by MBA can substantially enhance the execution capability and robustness of the robotic manipulation policies.
MBA accelerates the policy learning process, leading to more efficient learning (Q2). As shown in Fig. 2, we observe that policies with MBA exhibit higher learning efficiency during training compared to their vanilla counterparts. These policies typically reach the peak task test success rate at earlier training steps and maintain a more stable success rate at a higher level. This improvement is primarily attributed to the object pose information, which offers a more learnable and easily encoded feature representation.
V Real-World Experiments
In the real-world experiments, we aim to evaluate the effectiveness of the proposed MBA module in enhancing the performance of real-world robotic manipulation policies across a variety of manipulation tasks.
V-A Setup
Platform. We employ a Flexiv Rizon robotic arm with a Robotiq 2F-85 gripper to manipulate objects. A global Intel RealSense D415 RGB-D camera is positioned in front of the robot for 3D perception, capturing single-view workspace point clouds. The robot workspace is defined as a 40 cm $\times$ 60 cm rectangular area in front of the robot. The overview of the real-world robot platform is shown in Fig. 3. All devices are connected to a workstation with an Intel i9-10980XE CPU and an NVIDIA 2080 Ti GPU for data collection and evaluation. Since the training process of MBA requires 6D object motion data for supervision, we established a Motion Capture (MoCap) system comprising five OptiTrack Prime 13W infrared cameras. Before the training phase, we inpaint those MoCap markers within RGB-D observations to maintain visual consistency between the training and deployment. Notably, no motion capture is needed during deployment.
Tasks. We select three tasks for experimentation: Cut Clay (a tool-use contact-rich task), and Put Bread into Pot (soft body manipulation), and Open Drawer (articulated object manipulation), as illustrated in Fig. 3.
Demonstrations. We collect 50 expert demonstrations through end-effector teleoperation with haptic devices for each task, following the same setup and procedures in [rh20t, RISE, cage].
Baseline. We select RISE [RISE] as the baseline, which is a state-of-the-art real-world robotic manipulation policy that relies solely on 3D scene input and generates actions through a diffusion action head, conditioned on 3D features extracted via sparse convolutional networks [spatioenc] and transformers [transformer]. We then integrate the MBA module to predict object pose sequences prior to the action head of RISE, with additional conditioning on the generated object motion.
Protocols. In the real-world evaluation, we conduct 20 trials per method for each task, unless stated otherwise. All methods are compared under nearly identical randomized initial scene configurations for each trial.
V-B Cut Clay
The Cut Clay task consists of three stages: grasping the knife, slicing the clay until separation, and placing the knife on a foam mat. This task involves multiple stages, testing the ability of MBA to guide policy execution through accurate prediction of future object pose sequences.
In this experiment, we randomize the position and orientation of the cutting board, knife holder, and foam mat for each test. Special attention is given to varying the relative positions of these three objects to thoroughly assess policy robustness. Additionally, the shape of the clay is changed each time to meet generalization requirements for real-world applications. We define five evaluation metrics for this task:
  • Success rate of picking the knife.
  • Success rate of completing the cutting motion.
  • Success rate of slicing the clay until separation.
  • Success rate of placing the knife on the mat.
  • Total redundant knife-grasp attempts across 20 tests.
Methods Metrics
[…]
RISE $95$ $70$ $30$ $50$ $11$
RISE w. MBA $100$ $90$ $55$ $65$ $12$
TABLE III: Experimental results of the Cut Clay task.
From Tab. III, we observe that RISE with MBA shows significant improvements over RISE in all stages of task execution, particularly in the cutting and separation stage. This improvement is attributed to the need for rotating the cutting surface and splitting it when the knife cuts into the clay, which is inherently a 6-DoF task. MBA leverages the prediction of the tool motion to provide feedback on the execution of the current action, resulting in superior task performance compared to the baseline. However, MBA still exhibits repeated attempts to grasp the knife. We hypothesize that this issue arises from estimation errors in this task, where the thinness of the blade amplifies these errors, preventing them from being negligible.
Figure 3: Overview of our real-world robot platform and the illustration of three real-world tasks.
V-C Put Bread into Pot
The Put Bread into Pot task is a classic pick-and-place task, where the objective is to pick the bread from a bowl and place it into a pot. We chose this task because bread is a soft object. Unlike rigid bodies, soft objects can deform significantly due to the gripper’s hold and the pressure from touching the bottom of the pot, which affects the object’s pose, increasing the challenge for MBA in predicting the object pose sequence. We aim to assess the robustness and generalization ability of MBA through this task.
In this experiment, we randomly initialize the positions of the pot and the bowl and change the orientation of the bread within the bowl. We record two evaluation metrics: the average success rate of placing the bread into the pot and the total number of redundant attempts to grasp the bread.
Method Metrics
Success Rate (%) $\uparrow$ Attempt $\downarrow$
RISE $80$ $7$
RISE w. MBA $95$ $1$
TABLE IV: Experimental results of the Put Bread into Pot task.
The results, as shown in Tab. IV, indicate that RISE with MBA outperforms vanilla RISE by 15% in average success rate. In most cases, RISE with MBA successfully grasps the bread on the first attempt, whereas RISE often requires multiple attempts. This demonstrates that MBA can effectively handle object pose sequence prediction for soft objects, contributing to more accurate object localization and grasping.
V-D Open Drawer
The Open Drawer task is a two-stage process where the robot must first precisely grasp the drawer handle and then pull it open horizontally. The main difficulty lies in the small distance between the handle and the drawer surface, where even slight positional errors can result in a failed grasp or cause the gripper to lose contact during the pulling motion. This task challenges the policy’s ability to accurately estimate the handle pose sequence to ensure smooth and high-precision execution.
In this experiment, we account for the fact that delicate operations can be easily influenced by scene setup, particularly issues like point cloud occlusion when the robotic arm manipulates drawers at low heights. To minimize the impact of randomness, we doubled the number of test trials from the original plan, increasing it to 40 trials. Our test positions cover a variety of positions and orientations in the workspace, considering both fully closed and partially open states of the drawer. We also include scenarios where the initial drawer position causes point cloud occlusion, ensuring a comprehensive evaluation of the policies’ robustness.
We report the average success rates of the two task stages (Grasp: grasp the handle; Pull: pull the drawer until open) as evaluation metrics to accurately assess the execution capabilities of the policies.
Method Success Rates (%) $\uparrow$
Grasp Pull
RISE $37.5$ $37.5$
RISE w. MBA $52.5$ $52.5$
TABLE V: Experimental Results of the Open Drawer task.
The experimental results shown in Tab. V indicate that RISE with MBA outperforms the baseline RISE in both stages of the task. This aligns with our findings in simulation experiments, further confirming that integrating MBA significantly enhances the baseline policy in tasks that require precise object localization and fine manipulation.
VI Conclusion
In this paper, we introduce MBA, a novel module that draws inspiration from the reasoning process of human beings. MBA infers future object motion sequences and uses them as guidance for action generation. It can be flexibly integrated into existing robotic manipulation policies with diffusion action heads in a plug-and-play manner, significantly improving the performance of these policies across a wide range of tasks. This work holds great potential for future development, including integrating it into other robotic manipulation policies with different action heads, utilizing diverse object motion demonstration data (e.g., human demonstrations or web videos) for supervision and learning, exploring its performance in long-horizon, multi-stage tasks for both object pose and action prediction, and expanding it into a general large-scale policy across multi-dataset, multi-task settings.
[…]